Kate is applying to USA home loans for a mortgage of $200,000. The company is quoting 10.07% interest per annum. Kate would like to have a 25year amortization period and wants to make a monthly payment.

Required:

What will Kate’s payments be per month?

c) Exim Financial Services granted a loan of $22,000 to its client at 22% compound annual interest to be repaid over the next six (6)years. As a financial adviser to the client you are required to compute the size of the yearly payments.

d) A project costs $910,000 today and offers one million fifty thousand dollars cash flow in 72 months. It has a beta of 95 out of 100 and the risk premium on the market is 7.3%. if the treasury bill rate is 9%, should the project be accepted? Give reason.


1
Expert's answer
2020-11-24T05:07:30-0500

b) use the formula below 

"\frac{200000*(\frac{0,1007}{12})*(1+\frac{0,1007}{12})^{300}}{(1+\frac{0,1007}{12})^{300}-1}=1827.28"


c) find the formula:

"S=\frac{L\times i}{1-\frac{1}{(1+i)^n}}=\frac{22000\times 0.22}{1-\frac{1}{(1+0.22)^6}}=\frac{4840}{0.696721924}=6946.82"


d) find the discount rate

"R=Rf+\beta(Rm-Rf)=9+0.95(7.3-9)=7.385"

find NPV:


"NPV=-IC+\frac{CF}{(1+r)^n}=-910 000+\frac{1 050 000}{(1+0.07385)^6}=-910 000+684742.95=-225257.05"

NPV<0. NPV Is negative, therefore the project is unprofitable and money in such a project is not worth investing